A group of 4 students went to the movie theater. They each bought a ticket for $x and spent $10.50 each on snacks. A group of 3 adults went to the movie theater and paid twice as much for each of their tickets as each student. The group of adults also spent $10.00 each on snacks. The group of students spent the same amount as the group of adults. What is the cost of each adult ticket?
Found 4 solutions by greenestamps, Edwin McCravy, MathTherapy, ikleyn:
Answer by greenestamps(13200)



The wording of the problem is incorrect, or misleading. It says the students "each spent $10.50 on snacks", and then it says the adults "ALSO each spent $10.00 on snacks".

With the word ALSO in there, it seems those two amounts should be the same.

There is not much point in our trying to help you set up and/or solve the problem if we don't now for sure exactly what the problem is.


Answer by Edwin McCravy(20055)

A group of 4 students went to the movie theater.
They each bought a ticket for $x and spent $10.50 each on snacks.
They spent 4 times x dollars plus 4 times $10.50 or 4x+42 dollars. 

A group of 3 adults went to the movie theater and paid twice as much for each of
their tickets as each student, or $2x. The group of adults also spent $10.00
each on snacks.
They spent 3 times 2x dollars plus 3 times $10.00 or 6x+30 dollars. 

The group of students spent the same amount as the group of adults.
   3x+42 = 6x+30
      12 = 3x
       4 = x  <-- interpreted as $4.00 for student ticket.

What is the cost of each adult ticket?
Twice as much as $4.00 or $8.00.

Edwin

Answer by MathTherapy(10552)


A group of 4 students went to the movie theater. They each bought a ticket for $x and spent $10.50 each on snacks. A group of 3 adults went to the movie theater and paid twice as much for each of their tickets as each student. The group of adults also spent $10.00 each on snacks. The group of students spent the same amount as the group of adults. What is the cost of each adult ticket?
Cost of a student's ticket: $6
Cost of an adult's ticket: 2(6), or $12
Tutor EDWIN made a mistake. His equation should contain 4x + 42, instead of 3x + 42.

Answer by ikleyn(52781)

.

    4x + 4*10.50 = 3*(2x) + 3*10

    4x + 42      = 6x     + 30

    42 - 30      = 6x - 4x

    12           = 2x

     x           = 12/2 = 6.


Thus each student ticket price is 6 dollars.


ANSWER.  Each adult ticket price is 6*2 = 12 dollars.    ANSWER

Solved.